OpenStudy (anonymous):

"Laurence went to the store but forgot his wallet. He only has $20 to spend. He intended to buy 4lbs of cheese and 8lbs of sausage but that would cost $26. He decides to buy 3lbs of cheese and 5lbs of sausage which equals $17. How much do 2lbs of cheese and 4lbs of sausage cost?" This is my question and I already have "4c+8s=26, 3c+5s=17" but i have no idea what to do after that. And I also know that the answer is $13, i just have to figure out my work. "Laurence went to the store but forgot his wallet. He only has $20 to spend. He intended to buy 4lbs of cheese and 8lbs of sausage but that would cost $26. He decides to buy 3lbs of cheese and 5lbs of sausage which equals $17. How much do 2lbs of cheese and 4lbs of sausage cost?" This is my question and I already have "4c+8s=26, 3c+5s=17" but i have no idea what to do after that. And I also know that the answer is $13, i just have to figure out my work. @Mathematics

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4x + 8 = 26 3x + 5y = 17 understand?

OpenStudy (anonymous):

now you need to cancel out either x or y ...doesnt matter. I'm going to cancel out x first so multiply (4x + 8y = 26)3 (3x + 5y = 17)-4

OpenStudy (anonymous):

which gives you 12x + 24y = 78 -12x - 20y = -68

OpenStudy (anonymous):

so the x's will cancel out. leaving 4y = 10 and since you were looking for the price of 4 lbs of sausage... you know it cost 10$

OpenStudy (anonymous):

So can you find X and finish it up?

OpenStudy (anonymous):

?????? Are you there?

OpenStudy (anonymous):

Yes, I get it now...Thank you so much.

OpenStudy (anonymous):

No problem!!! :)